# Caution for the Environment: Multimodal Agents are Susceptible to Environmental Distractions

### v1

@@ Abstract
This paper investigates the faithfulness of multimodal large language model (MLLM) agents in the graphical user interface (GUI) environment, aiming to address the research question of whether multimodal GUI agents can be distracted by environmental context. A general setting is proposed where both the user and the agent are benign, and the environment, while not malicious, contains unrelated content. A wide range of MLLMs are evaluated as GUI agents using our simulated dataset, following three working patterns with different levels of perception. Experimental results reveal that even the most powerful models, whether generalist agents or specialist GUI agents, are susceptible to distractions. While recent studies predominantly focus on the helpfulness (i.e., action accuracy) of multimodal agents, our findings indicate that these agents are prone to environmental distractions, resulting in unfaithful behaviors. Furthermore, we switch to the adversarial perspective and implement environment injection, demonstrating that such unfaithfulness can be exploited, leading to unexpected risks. The code is available at https://github.com/xbmxb/EnvDistraction.
@@ 1 Introduction
Figure 1: (a) Previous studies expect agents to work normally and improve the action prediction performance (e.g., Yang et al. (2023)). (b) Recent works have discussed that agents can be distracted by user inputs through ambiguous instructions or malicious content (e.g.,Ruan et al. (2024)). (c) We focus on the setting where the agent is affected by distractions that exist in the environment. When the GUI agent is perceiving the environment, the screen presents distractions (e.g., coupons) that are irrelevant to the user’s goal. These distractions can interfere with the GUI agent’s task by distracting the agents from the user’s goal, thereby misleading the agent’s action trajectory.
@@
Empowered by the commendable progress in multimodal large language models (MLLMs) (OpenAI, 2023; Templeton et al., 2024), multimodal agents have demonstrated substantial potential in addressing complex interactive tasks Yao et al. (2022a); Shridhar et al. ; Wang et al. (2023). Graphical user interface (GUI) agents emerge as a prime example (Zhang and Zhang, 2023; Hong et al., 2023). GUI agents enable human-like behaviors on operating systems by receiving the environment status (e.g., screen) and predicting the subsequent action (e.g., “click the search bar”) based on a specific goal (e.g., “report hot financial news for today”). Their capabilities have reached an even more promising level with specialized augmentations. Concretely, related studies have confirmed that pre-planning and post-reflection are beneficial for interactive reasoning Hong et al. (2023); Zhang et al. (2024); Bai et al. (2024), and that layouts and localization are crucial for fine-grained perception Ma et al. (2024); Cheng et al. (2024); You et al. (2024). Benefiting from these studies, there is a growing trend of social impact to incorporate agents as AI assistants, thereby boosting efficiency and reducing human workloads in daily routines Wu et al. (2024b); Song et al. (2023).
@@
Despite the exciting progress, it remains an open question whether GUI agents can stay faithful to users without getting distracted Shi et al. (2023) by the rich contents in a multimodal environment. Figure 1-(c) shows a typical example. When deployed in the real-world community, GUI agents will inevitably face distractions, such as publicity and promotion activities, that interfere with their pursuit of user goals. If these distractions intervene in the agents’ actions, they may lead to uncontrollable environmental states, hindering agents from following the user’s intended goal. Even worse, the agents might complete an unexpected task suggested by the distractions.
On the one hand, those distractions cause significant changes in the action spaces for interaction. For example, the APP/website opening ads will cover the screen and disable all possible actions and some require manual shutdown, e.g., coupons. Under such a circumstance, the agents are likely to fall into an infinite loop if they stick to unavailable actions or fail to bypass the distractions. On the other hand, distractions also lead to inconsistency between the goal and the environmental contexts. The required operations are covered by distractions or mixed up with distractions, which means what exists is not (all) needed. This can be similar to conflicts in the inputs, where LLMs can be misled by unexpected content Mallen et al. (2023); Wei et al. (2023b); Shi et al. (2023); Li et al. (2023). As a result, agents will turn to another topic or even exhibit uncontrolled behaviors, instead of staying faithful to the original goal from the user instruction.
@@
This work focuses on faithfulness of multimodal GUI agents. Concretely, we explore the research question: To what extent can a GUI agent be distracted by a multimodal environment, thereby compromising its adherence to the goal? under the general circumstance where the user and the agent are both benign, the environment is risky but not malicious. Our research topic differs from existing work that either advances the GUI action performance or explores safety awareness but considers their comprehensive ability and reliability (Figure 1). Our considered distractions are less than ideal situations in an imperfect episode, such as interruptions, but are not abnormal adversarial examples, such as images with triggers.
@@
To address our research question, we first define the problem of distracting agents from the environment and three action types, gold, distracted, and invalid action. To simulate the distracting environment, we construct a dataset automatically, which contains four subsets, each for a vulnerable scenario, namely pop-up boxes, search, recommendation, and chat. Three working patterns are applied with different levels of perception and modal fusion. Experiments on 10 popular MLLMs reveal that both generalist and specialist GUI agents are susceptible to environmental distractions, and enhancing environmental perception is insufficient to mitigate these effects. To underscore the significance, we propose an adversarial setting, environment injection, and demonstrate that it is feasible to compromise an agent by leveraging such distractions.
@@
$\circ$ We construct a simulated dataset of distractions from the multimodal environment, empirically reveal the vulnerability of the agents’ faithfulness and present detailed analyses.
$\circ$ We demonstrate that the distractions can be exploited for malicious use with an automatic attacking named environment injection.
@@ 2 Related Work
This section introduces the background of MLLMs, GUI agents, and potential risks of language agents.
2.1 Multimodal Large Language Models
The remarkable capability of large language models has been successfully generalized to diverse modalities OpenAI (2023). A typical language-centered MLLM consists of a modality encoder, an LLM, an adapter (or a projector) as the bridge Liu et al. (2023); Alayrac et al. (2022); Zhu et al. (2023), and an optional generator for non-text modalities Wu et al. (2023); Shen et al. (2023). Recent works explore unified embedding Bavishi et al. (2023), grounding Peng et al. (2024), and encoder ensemble Hong et al. (2023), encouraging novel paradigms. Multi modalities enable the physical perception of agents, adding “eyes” to the language-based “brain” Zhang and Zhang (2023); Yan et al. (2023). The multimodal agents are facilitated to conduct actions in human-like behaviors.
@@ 2.2 Agents can Operate GUIs
Recently, the term agent has come to describe models that iteratively interact with an environment to solve complex problems Yao et al. (2022a, b). Among these challenges, GUI automation stands out as a representative task, demanding comprehensive perception and reliable actions.
Small-size models have achieved early success in GUI automation Sun et al. (2022); Rawles et al. (2023) by action selection. Since the emergence of LLMs Ouyang et al. (2022), the agents inherit language abilities from pre-training and the environment perception focuses on layout parsing or HTML code understanding Zhou et al. (2024); Lai et al. (2024). Empowered by multiple modalities, visual perception gradually replaces the complex strategies for the textual description of environments, advancing GUI agents to a generalizable paradigm. The visual channel plays a significant role in environment perception. Pre-training on element grounding Cheng et al. (2024), supplementing local layouts Ma et al. (2024), and selecting appropriate vision modules Zhang and Zhang (2023) have shown effective and essential in improving performance.
2.3 Risk of Language Agents
@@
$\circ$ The behavior of agents needs prejudgement. The behavioral risk lies in the implicit results of the agents’ actions. For example, agents should not forward unconfirmed gossip on social media. The risk is more concealed as it lies in the results rather than the generated content. Hence, detection and prevention require extrapolation based on prior knowledge and reasoning ability Tian et al. (2023); Yuan et al. (2024a). A representative work, Toolemu Ruan et al. (2024), emulates actions in a GPT-4-based sandbox.
Different from previous studies, our work proposes a novel setting (Figure 1) in three ways. (i) The possible distractions are in the environment instead of the user input. The distracting contents are received from the environmental perception. (ii) The user, agent, and environment are all benign, having no malicious intention or deliberate misleading. (iii) We focus on whether agents follow distracting content, instead of safety or ethics.
Figure 2: Overview of the methodology for distracting GUI agents. We first construct distracting environments (the left part), then implement working patterns with prompts (the middle part), and evaluate a broad range of multimodal agents, judging the predicted action as gold, distracted, and invalid (the right part).
@@ 3 Distracting GUI Agents
We begin by formalizing the problem statement in 3.1, then present approaches for distraction simulation in 3.2, measurement in 3.3, and working patterns 3.4. Figure 2 shows an overview.
@@ 3.1 Problem Statement
The formalization of GUI agent and the problem of distraction for GUI agents are as follows.
GUI agent. Consider a GUI agent $A$ interacting with an OS environment $Env$ to complete a specific goal $g$. At each time step $t$, the agent performs an action $a_{t}$ on the OS based on its perception of the environmental state $s_{t}$,
$\begin{split}&\textsc{Episode}=(g,[(s_{t},a_{t})]_{t=1}^{n}),\\ &a_{t}\leftarrow A_{LLM}(s_{t},g),s_{t+1}\leftarrow(s_{t},a_{t}),\end{split}$ (1)
@@
Distraction for GUI agents. However, the OS environment contains complex information of varying quality and from diverse sources. The environmental contents are formally divided into two categories: contents that are useful or necessary for goal completion, ${c^{use}}$, and distractors that are irrelevant to the user’s goal but indicate another target, ${c^{dist}}$,
@@
GUI agents must use $\{c_{t}^{use}\}$ for faithful actions instead of being distracted by ${c^{dist}}$ or predicting irrelevant actions. Based on the $s_{t}$, the available actions are determined. The valid action space $\mathbb{A}_{t}$ can be annotated with three types of labels, namely gold actions, distracted actions, and other actions,
$\begin{split}\mathbb{A}_{t}\leftarrow s_{t},\mathbb{A}_{t}=(\{a_{gold}\},\{a_{dist}\},\{a_{other}\}),\\ \end{split}$ (3)
where gold labels denote actions for $g$. Distracted labels denote actions for the distracting target, misled by distractions.
Receiving a goal $g$ and a screenshot as the current state $s_{t}$, the agent outputs the next action $a_{t}$. Then, $a_{t}$ is judged by comparing to action spaces. Our task concerns whether $a_{t}$ is faithful (gold) or distracted, or fails to be valid,
@@
The distraction problem needs to be studied under reasonable conditions, summarized as follows: First, be realistic. The environment state $s_{t}$, especially $c^{use}$, must reflect a realistic world. Second, be explicit. The distractors, $c_{dist}$, must be reasonable and distinguishable. It should be clear whether a piece of content is a distractor, similar to how humans can recognize ads. Third, be diverse. The environment states and distractors should exhibit diversity.
@@ 3.2 Distraction Simulation
Following the problem statement, we simulate the task without the loss of generality and construct a simulated dataset, $D$. Each sample is a triplet $(g,s,\mathbb{A})$ consisting of a goal $g$, a screenshot image representing the state $s$, and a valid action space $\mathbb{A}$. The critical part is to construct $s$ such that it includes $c^{use}$ and $c^{dist}$. We consider four common scenarios, namely pop-up boxes, search, recommendation, and chat, forming four subsets. A compositional strategy is employed for goals, layouts, and distractors. Table 1 shows an overview, including designs and sample numbers.
Scenario Pop-up boxes Search Recommendation Chat
@@
$\circ$ Pop-up boxes. Pop-up boxes frequently obscure the original web pages and prevent users from accessing the underlying content. This compels users to address the pop-up boxes before they can resume their intended actions, making the content of these boxes a significant distraction that must be managed. To simulate this scenario, we introduce pop-up boxes to a shopping website and modify their contents. Specifically, we leverage GPT-4 to generate target operations or questions as initial goals. For each goal, we then instruct GPT-4 to create various distractions, including common attractions like ads, notifications, and alerts. For layout, we prepare three templates for common pop-up boxes: (1) an input box with a submission button, (2) two buttons presenting choices, and (3) a four-choice checkbox. The faithful action is to dismiss the contents by clicking one of the buttons (such as “No thanks”) or by clicking a cross mark to close the box. If the agent follows the pop-up instead, it is considered distracted.
$\circ$ Search. Retrieval systems are found to be susceptible to the “Spiral of Silence” effect Chen et al. (2024), where LLM-generated (fake) text consistently outperforms human-authored content in search rankings, potentially leading to the marginalization of true information. Similarly, we examine the impact of integrating a fake item into search results. Concretely, we begin by prompting GPT-4 to generate common search queries. For each query, Google search API is called for the ground truth of the search results. Subsequently, we instruct GPT-4 to make up a distracting item, such as ads, different websites, or suggested other queries. The faithful action is to click on any of the true results. If the agent clicks on the fake results, it indicates a distraction from accurate information.
@@
$\circ$ Recommendation. The recommendation scenario presents related products according to the user query. We mix a fake product into the recommendation page for the shopping targets generated by GPT-4. For ground truth products, we simulate a BM25 Robertson et al. (2009) retriever on Amazon Reviews Hou et al. (2024). The product titles and descriptions are pre-indexed based on Lucene. The retrieval is limited to a specific category for efficiency. Similarly, GPT-4 makes up an appealing fake product to replace a random one. The action is considered faithful if the agent clicks any real product. This scenario differs from the search subset because the products are limited to the candidate set rather than the entire World Wide Web and the products are annotated with an accompanying picture.
@@
$\circ$ Chat. Chat is also a vulnerable scenario for GUI agents. Because the received messages will be displayed exactly as they are, which means a part of the screen can be decided by other information sources. We generate actions based on Discord webpage. Two goals with different actions are generated with the assistance of the Discord manual. One is rewritten to the user goal. The other is rewritten into a dialogue containing explicit action guides. The dialogues are posted to Discord by sending from two tool accounts. Then, the account that receives the action advice saves the chat record as a screenshot. The agent determines the next action for the user goal. If it follows the action guides in the dialogue, then it is distracted.
Action labels. The action labels are annotated during the generation. Given a screenshot, $\{a_{gold}\}$ and $\{a_{dist}\}$ are determined by the element layouts. Available actions that do not match the gold actions or follow the distractors are denoted as $\{a_{other}\}$, if any. $\{a_{inv}\}$ includes any actions that are not entailed in the layout. Based on the action labels, the targeting elements’ locations are annotated by OCR for coordinate evaluation for specialist agents.

### v2

# Caution for the Environment: Multimodal LLM Agents are Susceptible to Environmental Distractions

@@ Abstract
This paper investigates the faithfulness of multimodal large language model (MLLM) agents in a graphical user interface (GUI) environment, aiming to address the research question of whether multimodal GUI agents can be distracted by environmental context. A general scenario is proposed where both the user and the agent are benign, and the environment, while not malicious, contains unrelated contents. A wide range of MLLMs are evaluated as GUI agents using a simulated dataset, following three working patterns with different levels of perception. Experimental results reveal that even the most powerful models, whether generalist agents or specialist GUI agents, are susceptible to distractions. While recent studies predominantly focus on the helpfulness of agents, our findings first indicate that these agents are prone to environmental distractions. Furthermore, we implement an adversarial environment injection and analyze the approach to improve faithfulness, calling for a collective focus on this important topic. The code is available at https://github.com/xbmxb/EnvDistraction.
Caution for the Environment: Multimodal
LLM Agents are Susceptible to Environmental Distractions
Xinbei Ma1,2,3, Yiting Wang1, Yao Yao1,2,3, Tongxin Yuan1, Aston Zhang4, Zhuosheng Zhang1,∗, Hai Zhao1,2,3thanks: Corresponding authors. This research is supported by the Joint Research Project of Yangtze River Delta Science and Technology Innovation Community (No. 2022CSJGG1400), National Natural Science Foundation of China (62406188), and Natural Science Foundation of Shanghai (24ZR1440300). 1School of Computer Science, 2Key Laboratory of Shanghai Education Commission for Intelligent Interaction and Cognitive Engineering, Shanghai Jiao Tong University 3Shanghai Key Laboratory of Trusted Data Circulation and Governance in Web3 4GenAI, Meta {sjtumaxb, wyt_0416, yaoyao27, teenyuan, zhangzs}@sjtu.edu.cn, aston@meta.com, zhaohai@cs.sjtu.edu.cn
@@ 1 Introduction
Figure 1: (a) Previous studies expect agents to work normally and improve the action prediction performance (e.g., Yang et al. 2023, Zhang and Zhang 2023). (b) Recent works have discussed that agents can be influenced by ambiguous instructions or malicious inputs (e.g., Ruan et al. 2024). (c) We focus on the distractions from the environment. The agent is affected when it is perceiving the environment. These distractions (e.g., coupons) are irrelevant to the user’s goal and can mislead the agent’s action prediction.
@@
Empowered by the commendable progress in large language models (OpenAI, 2023; Templeton et al., 2024), agents have demonstrated significant potential in tackling interactive tasks Yao et al. (2022a); Shridhar et al. ; Wang et al. (2023), where GUI operating stands out as a prime multimodal example (Cheng et al., 2024; Hong et al., 2023). GUI agents replicate human-like behaviors on operating systems to achieve a specific goal (e.g., “report hot financial news for today”) by first understanding the environment status (e.g., screen) and then deciding the subsequent action (e.g., “click the search bar”). Their capabilities have reached an even more promising level through specialized augmentations: research has confirmed the value of pre-planning and post-reflection for overall trajectories Hong et al. (2023); Zhang et al. (2024a), as well as the importance of localized layout grounding for perception. Ma et al. (2024); Cheng et al. (2024); You et al. (2024). Building on these studies, there is a growing societal trend to adopt AI agents as assistants, boosting efficiency and alleviating human workloads Wu et al. (2024c); Song et al. (2023).
@@
Despite the exciting progress, it remains an open question whether GUI agents can stay faithful to user intentions without getting distracted Shi et al. (2023) by the rich contents in the environment. Figure 1-(c) shows a typical example. When operating in real-world scenarios, GUI agents are inevitably exposed to distractions that can interfere with their pursuit of user goals, such as publicity and promotion activities. If these distractions influence the agents’ actions, they may lead to uncontrollable environmental states. Even more concerning, the agents might complete an unexpected task suggested by the distractions.
@@
This work focuses on the faithfulness of multimodal GUI agents. Concretely, we explore the research question: To what extent can a GUI agent be distracted by a multimodal environment, thereby compromising its adherence to the goal? under the general circumstance where the user and the agent are both benign, the environment is risky but not malicious. As illustrated in Figure 1, our study differs from existing work that either advances the GUI action performance or explores safety awareness. We consider general, imperfect situations, neither assuming an ideal environment nor simulating abnormal adversarial attack situations.
@@
Our study begins with defining the problem of environmental distraction for GUI agents. We construct a dataset comprising four subsets, each designed to simulate a vulnerable scenario involving distractions: pop-up box, search, recommendation, and chat. We then propose three working patterns that differ in their levels of perception and modality fusion. Experiments on ten popular MLLMs reveal that both generalist and specialist GUI agents are susceptible to environmental distractions. Furthermore, simply enhancing environmental perception proves insufficient to mitigate this lack of faithfulness. In the analysis, we introduce a faithfulness improvement method by adding preference to the inputs. Finally, we implement adversarial environment injection, demonstrating the feasibility of compromising an agent through these distractions.
@@
$\circ$ We construct a simulated dataset of distractions from the multimodal environment, empirically reveal the vulnerability of the agents’ faithfulness, and present detailed analyses.
$\circ$ We analyze the malicious use of distractions for environment injection and the improvement approach for faithfulness.
@@ 2.1 Agents can Operate GUIs
Recently, the term “agent” has been used to refer to models that interact with an environment to solve complex tasks Yao et al. (2022a, b). Among these challenges, GUI automation stands out as a representative task, demanding comprehensive perception and action prediction.
Small models have achieved early success in action selection Sun et al. (2022); Rawles et al. (2023). Since the emergence of LLMs Ouyang et al. (2022), the agents inherit language abilities and interpret the environment by HTML code understanding Zhou et al. (2024); Lai et al. (2024). Empowered by multimodal pre-training, visual perception gradually replaces the textual description of environments, allowing GUI agents to look at the screen. Hence, visual augmentation plays a significant role in environment modeling and performance improvement Cheng et al. (2024); Ma et al. (2024); You et al. (2024).
Figure 2: Overview of our work for distracting GUI agents. We first construct environment status with distractions (the left part), then implement working patterns with prompts (the middle part), and evaluate a broad range of multimodal agents, judging the predicted action as gold, distracted, and invalid (the right part).
2.2 Potential Risk of Agents
@@
$\circ$ The behavior of agents needs prejudgement. The risk is more concealed as it lies in the implicit results rather than the literal meaning Liao et al. (2024); Zhang et al. (2024b). Hence, detection and prevention require extrapolation Tian et al. (2023); Yuan et al. (2024a); Hua et al. (2024); Zhang et al. (2024c). A representative work, Toolemu Ruan et al. (2024), emulates actions in a GPT-4-based sandbox.
Different from previous studies, our work proposes a novel setting (Figure 1) because (i) The distractions are received from the environment instead of malicious input. (ii) All roles are benign without malicious intention or deliberate misleading. (iii) We focus on whether agents follow distractions, instead of safety or ethics. We aim to reveal a general unfaithful risk rather than carefully crafted adversarial attacks.
@@ 3 Distracting GUI Agents
We begin with the problem statement in Section 3.1, then introduce approaches for distraction simulation in Section 3.2, measurement in Section 3.3, and working patterns in Section 3.4. Figure 2 shows an overview.
@@ 3.1 Problem Statement
GUI agent. Consider a GUI agent $A$ interacting with an OS environment $Env$ to complete a specific goal $g$. At each time step $t$, the agent perceives and understands the environmental state $s_{t}$ and decides an action $a_{t}$ to perform on the OS,
$\begin{split}&a_{t}\leftarrow A_{LLM}(s_{t},g),s_{t+1}\leftarrow(s_{t},a_{t}),\end{split}$ (1)
@@
Distraction for GUI agents. The environment contains complex information of varying quality and from diverse sources, formally divided into two parts: contents that are useful or necessary for achieving the goal, ${c^{use}}$, and distractions that are irrelevant to the user’s goal and may suggest another target, ${c^{dist}}$,
@@
The valid action space $\mathbb{A}_{t}$ is determined by $s_{t}$ and can be annotated with three types of labels, i.e., gold actions, distracted actions, and other actions,
$\begin{split}\mathbb{A}_{t}\leftarrow s_{t},\mathbb{A}_{t}=(\{a_{gold}\},\{a_{dist}\},\{a_{other}\}).\end{split}$ (3)
GUI agents must use $\{c_{t}^{use}\}$ to predict a gold action instead of following ${c^{dist}}$ to predict a distracted action or generate other irrelevant actions. By comparing to the labeled action space, $a_{t}$ is judged to be faithful (gold), distracted or fails to be valid,
@@ 3.2 Distraction Simulation
Following the problem statement, we construct a simulated dataset, $D$. Each sample is a triplet $(g,s,\mathbb{A})$ consisting of a goal $g$, a screenshot image as environment state $s$, and a valid action space $\mathbb{A}$. Since existing datasets cannot be used directly, our core idea is to make a realistic screenshot suitable for our task with minimal modification by inserting a realistic distraction. Specifically, the simulation of distraction is carefully decomposed into detailed steps, resulting in a compositional strategy for layouts, goals, and distractions. Algorithm 1 presents the unified pipeline of data construction, followed by the descriptions of four subsets, each for a common scenario, namely Pop-up box, Search, Recommendation, and Chat. The final overview and statistics are shown in Table 1.
Algorithm 1 Distraction simulation
1:  Initialize: Website template $s_{template}$, Target layouts $S_{target}$, LLM, external tool $T$, Maximum tries $t_{m}$.
2:  Notions: User’s goal $g$, Distracting goal $d$, action space $\mathbb{A}$.
3:  for $\{s_{target}\}\in S_{target}$ do
4:     for $t<t_{m}$ do
5:        $g\leftarrow\text{LLM}(s)$,
6:        $d\leftarrow\text{LLM}(s),d\neq g$
7:        $c^{use}\leftarrow\text{LLM}(s_{target},g,T)$
8:        $c^{dist}\leftarrow\text{LLM}(s_{target},d)$
9:        $\mathbb{A}$ is determined by $c^{use}$ and $c^{dist}$
10:        $s_{target}^{\prime}\leftarrow s_{target}+c^{use}+c^{dist}$
11:        $s_{template}\leftarrow s_{template}+s_{target}^{\prime}$
12:        $t\leftarrow t+1$
13:     end for
14:  end for
Pop-up box Search Recommendation Chat
@@
$\circ$ Pop-up box. The initial template is a homepage of a webshop written in HTML, and we prepare three templates of common pop-up boxes for target layouts (Line1): one submission button, two options, and a four-option checkbox. The faithful action is to dismiss the contents by clicking one of the buttons (such as “No thanks”) or by clicking a cross mark to close the box. If the agent follows the pop-up instead, it is considered distracted. We prompt GPT-4 to generate initial goals (Line5). For each goal, GPT-4 creates various distractions including ads, notifications, and alerts (Line6). After filled with headlines and button names (Line7-8), the popup box is inserted into the homepage, displayed in the browser and the screenshot is taken (Line11).
$\circ$ Search. AI-generated contents are found to undermine retrieval systems by marginalizing true information Chen et al. (2024). This subset simulates the impact of inserting a fake result into search results, based on the template layout of the search result webpage. We generate common search queries (Line5) and call Google Search API to retrieve the real search results for each query (Line7). Subsequently, distracting results generated by GPT-4 are inserted (Line8-11). The faithful action is to click on any of the true results. If the agent clicks on the fake results, it indicates a distraction from accurate information.
@@
$\circ$ Recommendation. The recommendation webpage presents related products according to the user query. We follow a product display webpage as the target layout and mix an AI-generated product into the recommended products for each shopping target. Unlike the worldwide search engine, our recommendation system simulates a BM25 Robertson et al. (2009) retriever on Amazon Reviews Hou et al. (2024) (Line7). Similarly, GPT-4 makes up an appealing fake product to replace a random one. This scenario differs from the search subset because of the quality of real results. The product retriever is constrained by the limitations of the candidate set, while the search engine accesses the entire World Wide Web.
@@
$\circ$ Chat. In a chat window, received messages are displayed exactly as sent, meaning that a portion of the screen is controlled by external information sources. This subset leverages the Discord chat room. Two different goals are generated based on the Discord manual (Line5-6). One is rewritten to the user’s goal, and the other is rewritten into a dialogue providing explicit action guides as the distraction (Line7-8). The dialogues are posted to the chat server from two tool accounts, shown on the screen (Line11). The agent determines the next action for the user goal. If it follows the action guides in the dialogue, then it is distracted.
Action labels. During the above process, $\{a_{gold}\}$ and $\{a_{dist}\}$ are determined by $c^{use}$ and $c^{dist}$. Other possible actions are labeled as $\{a_{other}\}$, if any. Related locations on the screenshots are annotated by OCR to evaluate the coordinate prediction of specialist agents.